OpenStudy (anonymous):

(the v means downards, these 3's are subscripted) solve for x: logv3(x-8) + logv3(x) = 2 Please be specific! Thank you! :)

OpenStudy (unklerhaukus):

underscore means subscript

OpenStudy (unklerhaukus):

\[\log_3(x-8) + \log_3(x) = 2\] \[=\log_3((x-8)x)=2\]

OpenStudy (unklerhaukus):

now use the definition of logarithm \[\log_bx=y\qquad\rightarrow\qquad b^y=x\]

OpenStudy (unklerhaukus):

\[3^2=(x-8)x\]

OpenStudy (anonymous):

Since this is a logarithm is is it 2x-8x = 9 or x^2-8x = 9? Is the final answer = 2/3? If not, how do I solve x^2-8x = 9?

OpenStudy (unklerhaukus):

\[x^2-8x = 9\]\[x^2-8x -9=0\]\[=(x+...)(x-...)\]

OpenStudy (unklerhaukus):

looking for two numbers that have a difference of 8 and a product of -9

OpenStudy (unklerhaukus):

now there product is negative so one of the numbers must be negative

OpenStudy (anonymous):

x = 9 x = -1right?

OpenStudy (anonymous):

x = -1 is extraneous?

OpenStudy (unklerhaukus):

\[x^2−8x−9=0\] \[(x+1)(x-9)=0\] \[x=-1,9\]

OpenStudy (anonymous):

x=9 in first part and x=-1 in second is illegal

OpenStudy (anonymous):

in real log we can not chose negative variable

OpenStudy (unklerhaukus):

\[\log_3(x-8) + \log_3(x) = 2\] if we substitute x=-1 in to original equation \[\log_3(-1-8) + \log_3(-1) = 2\] \[\log_3(-9) + \log_3(-1) = 2\] but you cannot take the log of a negative number if we substitute x=9 in to original equation \[\log_3(9-8) + \log_3(9) = 2\]\[\log_3(1) + \log_3(9) = 2\]\[\log_3(9)=2\]\[3^2=9\]which is true

OpenStudy (anonymous):

ok i guess it was 8-x in log so only x=9 is answer

OpenStudy (anonymous):

Thank you so much to both of you guys!! This helped a lot :)